Circular No. 9095

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 W1 (HILL) R. E. Hill, Lunar and Planetary Laboratory, reports his discovery of a

comet (discovery observation tabulated below) on CCD images taken with the 0.68-m Schmidt telescope in the course of the Catalina Sky Survey; three co-added 30-s exposures in 4′′–5′′ seeing show a small nuclear condensation 7′′–8′′ across with a diffuse, faint tail 20′′–30′′ long in p.a. ∼ 300^{o}. Four coadded 40-s exposures in 2′′–3′′ seeing by Hill on Nov. 19.4 UT show a nuclear condensation 4′′–5′′ across with a 1′′–2′′ coma and a tail 20′′–30′′ long and 10′′–12′′ wide in p.a. ∼ 300^{o}. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD astrometrists have also noted the object’s cometary appearance. F. Fratev, Sofia, Bulgaria, reports that twenty-one stacked 40-s CCD exposures obtained with a 0.35-m reflector at Plana, Bulgaria, on Nov. 19.1 shows a diffuse coma nearly 10′′ in diameter and a tail ∼ 60′′ long in p.a. 275^{o}. G. Hug (Scranton, KS, U.S.A., 0.56-m reflector, Nov. 19.5) notes a condensed nuclear condensation and a short, diffuse tail 20′′ long in p.a. 290^{o}. A. R. Gibbs (Mt. Lemmon 1.5-m reflector, Nov. 19.5) writes that four co-added 60-s unfiltered CCD images taken in good seeing show a bright, diffuse coma of size 5′′ × 7′′ with a 30′′ tail in p.a. 290^{o} whose width is constant and equal to the coma diameter. D. Chestnov, Moscow, reports that five stacked 300-s unfiltered CCD exposures taken by T. Kryachko and B. Satovski with a 0.3-m reflector at the Zelenchukskaya Station of Engelhardt Observatory on Nov. 19.04 show a tiny coma and a 0′.2 tail in p.a. 285^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
Nov. 18.50653 9^{h}46^{m}17.^{s}11 +28^{o}51′20′′.0 18.5 Hill

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-W34; the true orbit may well be of short period.

T = 2009 June 26.522 TT  ω = 18.^{o}740
Ω = 54.303 }2000.0
q = 2.46311 AU  i = 14.413

V496 SCUTI Visual magnitude estimates (cf. IAUC 9093): Nov. 14.724 UT, 7.6 (P.

Schmeer, Bischmisheim, Germany); 18.719, 7.2 (Schmeer); 18.722, 7.3 (K. Hornoch, Vohancice, Czech Republic).

2009 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green